OpenStudy (anonymous):

slope=1/2, passes through (6,4) i have NO idea what im doing.

OpenStudy (anonymous):

here is the equation we will be using: y=mx=b okay so were trying to find B to find b you substitite y with 4 and x with 6 m=slope and the solpe is 1/2 so y=mx=b 4=1/2(6)+b multiply 1/2times 6 so, 4=3+b now we have to get that b alone! so you subtract the 3 from both sides because what you do to one side you must do to the other! 3 crosses out on the right so b =1 :D hope that helped!

OpenStudy (anonymous):

whoopse i mean y= mx +b